CALCULATED SPECTROSCOPIC PROPERTIES FOR CLUSTER ANIONS OF TYPE Hal^{−} · ··HCCH (Hal: F, Cl, Br, and I)

\underline{P. BOTSCHWINA}, R. OSWALD, S. SCHMATZ, Institut für Physikalische Chemie, Universität Göttingen, Tammannstraße 6, D-37077 Göttingen, Germany; and H. STOLL, Institut für Theoretische Chemie, Universität Stuttgart, Pfaffenwaldring 55, D-70569 Stuttgart, Germany.

Large-scale coupled cluster calculations were carried out for the complexes formed by a halide anion and an acetylene molecule, all of which have linear equilibrium structures ^{a} ^{b}. The complex [F⋯H⋯CCH]^{−} is most strongly bound and extremely anharmonic. A comparative discussion of various spectroscopic properties is given.

aP. Botschwina, T. Dutoi, M. Mladenović, Rainer Oswald, S. Schmatz and H. Stoll, Faraday Discuss., in press.

bP. Botschwina and H. Stoll, Phys. Chem. Chem. Phys., in press.